Question 3

An automated egg carton loader has a 3% probability of cracking an egg, and a customer will complain if 1 or more eggs are cracked. Assume each egg load is an independent event. What is the probability that a box of a 12 eggs results in a complaint?

a.

o.25

b.

0.31

c.

0.47

d.

0.73
Answer #1

Answer

We know that the probability is given as p= 3% = 3/100 = 0.03

Sample size is given as n = 12

There will be a complaint only if 1 or more eggs are cracked.

So, we can say that

Probability of one or more eggs cracked = 1- Probability of zero egg cracked. .............................equation 1

Using the binomial formula

P(x=r) = inom{n}{r}*p^r*(1-p)^{n-r}

where r = 0, n = 12 and p = 0.03

setting the values, we get

P(z = 0) = ( ) * 0.030* (1-0.03)12-0 = (121/((12-0)! * 0!)) * 1 * 0.9712 12!/((12 -0)! 0!) 1 0.97

this gives us

Plz. = 0)-( 12y( (12-0)! * 0!)) * 1 * 0.9712-0.6938

setting this value in equation 1, we get

Probability of one or more eggs cracked = 1- Probability of zero egg cracked = 1 - 0.6938 = 0.3062 or 0.31

So, option B is correct answer